# LLM-SQL-Solver: Can LLMs Determine SQL Equivalence?

### v2

@@ Abstract
Judging the equivalence between two SQL queries is a fundamental problem with many practical applications in data management and SQL generation (i.e, evaluating the quality of generated SQL queries in text-to-SQL task). While the research community has reasoned about SQL equivalence for decades, it poses considerable difficulties and no complete solutions exist. Recently, Large Language Models (LLMs) have shown strong reasoning capability in conversation, question answering and solving mathematics challenges. In this paper, we study if LLMs can be used to determine the equivalence between SQL queries under two notions of SQL equivalence (semantic equivalence and relaxed equivalence). To assist LLMs in generating high quality responses, we present two prompting techniques: Miniature & Mull and Explain & Compare. The former technique is used to evaluate the semantic equivalence in which it asks LLMs to execute a query on a simple database instance and then explore if a counterexample exists by modifying the database. The latter technique is used to evaluate the relaxed equivalence in which it asks LLMs to explain the queries and then compare if they contain significant logical differences. Our experiments demonstrate using our techniques, LLMs is a promising tool to help data engineers in writing semantically equivalent SQL queries, however challenges still persist, and is a better metric for evaluating SQL generation than the popular execution accuracy.
@@ 1 Introduction
Structured Query Language (SQL) it is the standard programming language to query data systems and it is one of the most popular programming language  [^1] https://www.stackscale.com/blog/most-popular-programming-languages/. Since SQL is not a procedural language, one may write different SQL queries that are semantically equivalent. As a result, determining the equivalence of SQL queries is a fundamental task in many practical important applications ranging from query optimizers in database systems Pirahesh et al. (1992); Bai et al. (2023); Zhou et al. (2021); Samwel et al. (2018) to evaluating the translation from user text-to-SQL Zelle and Mooney (1996); Zhong et al. (2020b); Khatry et al. (2023); Floratou et al. (2024). However, it has been shown by the data management research community that checking the equivalence of two SQL queries is in general undecidable Abiteboul et al. (1995). While there are theoretical works to check equivalence under different assumptions Chandra and Merlin (1977); Sagiv and Yannakakis (1980); Cohen et al. (1999); Chu et al. (2018) and implemented prototypes such as Cosette Chu et al. (2017), they only support limited SQL operations and such projects are no longer maintained due to lack of experts. For instance, Cosette does not support commonly found operations like EXISTS, IN, queries with constant expressions, etc.; and Cosette is no longer maintained due to the lack of staff. Query optimizers and query processing engines in data systems still primarily rely on heuristic query rewriting rules Begoli et al. (2018). Exploring and writing these rules correctly require strong domain expertise and need many engineering hours.
@@
Moreover, due to the difficulty in evaluating the equivalence of two SQL queries, Text-to-SQL literature Zhong et al. (2017); Yu et al. (2018); Wang et al. (2019); Lee et al. (2021); Li et al. (2023) primarily use exact match (exact SQL string match) or execution match (compare the execution outputs) as the evaluation metrics Zhong et al. (2020a). However, exact string match may overlook many equivalent SQL queries (there are many ways different ways to write the same SQL query) and execution match may misclassify different SQL queries as equivalent due to no counterexamples are contained in the underlining database instance.
@@
Recent advancements in large language models (LLMs) have shown great capability in following instructions and in generating decisions that align with human judgement Bubeck et al. (2023); Chiang and Lee (2023); Yan et al. (2023a); Zhang et al. (2023b). For instance, It has been shown that GPT models are capable of performing first order logic reasoning  Saparov and He (2022), detect implicit contents Huang et al. (2023); Zhang et al. (2023c), and refining the invisible logic Feng et al. (2023). In this paper, we investigate whether LLM can act as a judge to determine the equivalence of SQL queries. Hence, in this paper, we would like to mainly answer the following questions:
@@
  • Should LLMs be used to evaluate the quality of generated SQL queries with the expected SQL answer?
@@ 2 Background
2.1 Motivation
To demonstrate the capability of LLMs, we first conduct an initial study on a real-world query rewrite error which is known as the COUNT bug Ganski and Wong (1987). Here are two SQL queries Q1 and Q2 as shown below:
@@ 11  FROM Supply
At first sight, one may think these two queries (the COUNT bug) are equivalent, since they both return results based on the condition that qoh from Supply table equals to the counts of records in the Supply table when shipdate less than 10; however they are different in how they handle the subquery. Suppose that there is a record $R$ in the Parts table such that R.qoh is zero and R.pnum does not exist in the Supply table. Then, Q1 will include record $R$ since R.qoh is zero and selecting the count for R.pnum from the Supply table will also return zero. On the other hand, Q2 will not include record $R$ because the TEMP result will not contain R.pnum.
We gave this equivalence question to ChatGPT and ChatGPT is able to correctly identify that these two SQL queries are not equivalent. When asked to provide a counterexample, it is able to generate a counterexample in which it provided a database instance containing such a record $R$ as described above.
While proving the semantic equivalency between two SQL queries has been proved to be undecidable Trahtenbrot (1963); Libkin (2004); Sagiv and Yannakakis (1980), we believe LLMs may offer new opportunities to help data engineers in scrutinizing their SQL optimization logic and potentially identify flaws accordingly.
@@ 2.1.1 Two Notions of SQL Equivalence
First, we define two different notions on SQL equivalence motivated from query optimizers in data systems and evaluation from the text-to-SQL literature.
@@
Based on the semantic definition of SQL equivalence (Definition 2.1), the Count bug is indeed a flaw for database’ query engines and should be remedied immediately. On the other hand, we argue that based on Definition 2.2, which reflects the logical perspective and aligned with human perception, the two SQL queries in the Count Bug example is relaxed logically equivalent. The fact that even experienced experts failed to identify this subtle difference indicates humans often recognize equivalence based on the relaxed notion. As a result, we believe that the relaxed notion of SQL equivalence can better capture a human user’s satisfaction with the generated SQL from the foundational models. We asked ChatGPT if the two queries in the Count Bug are relaxed equivalent and it successfully identified them as relaxed equivalent because they share the same logic.
To illustrate why these two queries are relaxed logically equivalent, Q2 in the Count Bug example only the TEMP table needs to be edited as shown below for it to be semantically equivalent to Q1 in the Count Bug example:
1Q2 TEMP Table Original:
2WITH Temp AS
3  SELECT Supply.pnum, COUNT(Supply.shipdate) AS ct
4  FROM Supply
5  WHERE Supply.shipdate < 10
6  GROUP BY Supply.pnum
7Q2 TEMP Table After Edits:
8WITH Temp AS
9  SELECT Parts.pnum, COUNT(Supply.shipdate) AS ct
10  FROM Parts, Supply
11  WHERE Supply.shipdate < 10
12  AND Parts.pnum = Supply.pnum
@@ 10  FROM Parts, Supply
For text-to-SQL task, we believe that when comparing the gold SQL query (expected answer) with the generated SQL query, the current most popular metric, execution match, does not reflects the human perspective and fails to capture a user’s opinion on the generated SQL. For instance, a common failure reasons in execution match is that the generated SQL selects fewer/extra columns than the gold SQL. In fact, a recent work has explicitly added "do not select extra columns" in its prompt to improve the execution accuracy Dong et al. (2023). However, from user’s perspective providing extra columns may not reduce the quality of the answer. Also, there are many other similar reasons that can lead to false execution matches but not necessarily indicate a lower quality of the generated SQL query (such as the column order in selection, case sensitivity, and duplicate entries or extra aggregates in the output). Furthermore, execution matching fails to correctly identify semantically different queries when the underlying database does not contain any counter examples, even when developer spend hundreds of hours on creating the benchmark Yu et al. (2018). As a result, it would be desirable to have a SQL solver which can determine SQL equivalence without constructing sophisticated databases and without accessing any stored data. Hence, in additional to investigating the capability of foundational models in identifying semantic equivalence, we further explore whether a foundational model can be used as a better metric for text-to-SQL generation benchmarks.
@@ 3 Proposed Method
For semantic equivalence, we also added a short definition of SQL semantic equivalence (i.e., Note: Two SQL queries are semantically equivalent if and only if they yield identical results for all possible databases.). In addition, we first ask GPT to generate a counter example if there exists one and then decide if two SQL queries are equivalent. We call this prompt the Counter Example prompt. However, as shown in Figure 1, the counter example prompt leads to sub-optimal results, especially when GPT models only scratched the surface and observed the obvious (similarity between two SQL queries in structure and logic). This finding aligns with the prior works Wang et al. (2023); Zheng et al. (2023) such that GPT models exhibit bias toward ignoring small differences when determining SQL equivalence and exhibit limit reasoning capability.
To overcome the observed phenomena, we introduce our proposed prompting technique in detail to obtain better responses from large language models. In complex problem solving, people always start with a simple example and then introduce more complex elements to understand the full picture. Motivated by this simple intuition, we incorporate this idea and proposed the Miniature and Mull prompting technique. In particular, we first ask GPT to execute both queries on a simple database and then see if it can modify the database such that the two queries output different results (i.e., a counter example). As shown later in the experimental section, this approach can directly enhance the LLM’s response accuracy in determining SQL semantic equivalence. We believe that this prompting technique has wide applications in comparing the equivalence of two codes, plans, and logic.
For relaxed equivalence, we designed the Explain and Compare prompt by asking the LLMs to perform multi-step reasoning before determining if the two SQL queries are equivalent or not. This prompt follows the main ideas from the chain of thoughts Wei et al. (2022). We first prompt the LLMs to explain two SQL queries and then ask it to determine if these two queries contain significant difference in logic, as shown in Figure 1. We observe that LLMs are inclined to label two SQL queries are equivalent when they appear similar in structure. As a result, it is crucial for LLMs to be able to first spot the differences between these two queries and then apply this knowledge in considering if these two SQL queries are relaxed equivalent or not.
@@ 4 Evaluations
We evaluate the capability of popular large language models in identifying semantic and relaxed equivalence between a pair of SQL queries. In particular, we showcase the experimental results using the popular GPT 3.5 turbo (gpt-3.5-turbo) and GPT 4 turbo (gpt-4-1106-preview) models. The models are accessed via OpenAI API.
Hyper-parameter: We use zero demonstrations in our prompts. In addition, in the API call, we have set temperature to 0 which implies greedy sampling is used to generate output; The max tokens are set to 3000; and the stopping tokens are set to "Answer:" for all experiments.
@@ 4.1 Data Source
Our evaluation was conducted on i) Non-equivalent Queries Chu et al. (2017) and ii) Spider cross-domain benchmark Yu et al. (2018). The Nonequivalent Queries are originally used to evaluate the performance of Cosette a formal language based SQL solver; however, Cosette does not support many SQL operations such as IN, EXCEPT, etc.; and the Cosette project is no longer maintained. We use Nonequivalent Queries to investigate if LLMs can successfully notice the semantic differences (often subtle) between two non-equivalent SQL queries.
  • Nonequivalent Queries Chu et al. (2017): It contains 4 challenging and 1 easy questions, including the COUNT BUG, one exam question from CSE344 (data management) at the University of Washington, two bugs users found on SQL solver, and a simple hello world example.
The Spider benchmark contains 1034 development questions covering 20 databases in different application domains, and each database contains multiple tables. The Spider SQL queries are classified into four levels of complexity: easy, medium, hard, and extra. This classification is based on various factors, including the quantity of SQL keywords utilized, the incorporation of nested subqueries, and the usage of column selections and aggregations.
The official evaluation metrics on the Spider benchmark are execution accuracy and exact string match accuracy. Since two semantically equivalent SQL queries can be written in many different ways, execution accuracy is often consider as more important than exact string match. In all experiments, we use the open sourced execution match evaluation code from GitHub  [^2] https://github.com/THU-BPM/ChatGPT-sql. However, as we argued in previous sections, execution match neither captured the semantic equivalence notion nor the relaxed equivalence notion. As a result, we showcase the results of using LLMs to determine if the generated SQL queries are the equivalent or not. We used three open-sourced results.
  • Chat Liu et al. (2023): Generate SQL queries using ChatGPT with zero-shot basic prompt.
  • DIN Pourreza and Rafiei (2023): Generate SQL queries by decomposing the task into fine grained sub-modules and use the GPT-4 model with few-shot prompting
@@
4.2 Finding Counter Examples
Judge Prompt Accuracy
GPT 3.5 turbo CounterExample 4/5
Miniature&Mull 5/5 $\uparrow$
GPT 4 turbo CounterExample 3/5
Miniature&Mull 5/5 $\uparrow$
GPT 4 CounterExample 4/5
Miniature&Mull 4/5 -
Table 1: Results in Nonequivalent Queries
We test LLMs ability in deciding if two SQL queries are semantically equivalent. Recall, two queries can not be semantically equivalent if there exists a counter example (a database instance) such that two queries produce different outputs.
We used two different prompting methods Counter Example and Miniature & Mull to compare and contrast the accuracy of LLMs. An answer is accurate if the answer states the two SQL queries are not equivalent.
Judge Data Source Equivalence/Prompt Easy Medium Hard Extra Acc
@@ 4.2 Finding Counter Examples
Table 2: A breakdown of the accuracy scores when using different metrics, i.e., execution accuracy and LLMs judgements with varying prompt techniques. The compared SQL queries are the gold SQL queries from Spider and the generated SQL queries from Chat, DIN, and DAIL.
As shown in Table 1, when using the Counter Example prompt (asking LLMs to determine SQL equivalence and provide a counter example if two SQL queries is not equivalent), GPT 4 and GPT 4 turbo failed to notice the subtle difference in the COUNT BUG, whereas GPT 3.5 turbo correctly output that queries in COUNT BUG are not semantically equivalent. In addition, GPT 4 turbo failed on a user found bug originally posted on Hackers News  [^3] https://github.com/uwdb/Cosette/issues/29. This bug contains two SQL queries where one used DISTINCT in selection and the other used UNION ALL to union two sub-queries. However, GPT 4 turbo failed to recognize that UNION ALL will not remove duplicate rows (though UNION will). GPT 3.5 turbo failed on the CSE344 exam question which involves different sub-queries.
When using Miniature & Mull (asking LLMs to first run a simple example over the given SQL queries and then examining if the example database can be modified to be a counter example), GPT 3.5 turbo and GPT 4 turbo successfully identified all the mistakes. However, GPT 4 still failed on the COUNT BUG. We can observe that, when using Counter Example prompt (follows the main ideas of chain of thoughts), LLMs scores at least 60% accuracy. When using the Miniature & Mull prompting, the accuracy is significantly improved to 100% for GPT 3.5 turbo and GPT 4 turbo. As a result, we believe that letting LLMs execute the SQL queries on a simple database instance and then considering a counter example is an effective way to determine SQL semantic equivalence. Moreover, since GPT 4 turbo showcases strong performance with faster response time and less monetary costs [^4] https://openai.com/pricing compared to GPT 4, we use GPT 3.5 turbo and GPT 4 turbo for all the following experiments.

### v3

@@ Abstract
Judging the equivalence between two SQL queries is a fundamental problem with many practical applications in data management and text-to-SQL generation. While the research community has reasoned about SQL equivalence for decades, it poses considerable difficulties and no complete solutions exist. Recently, Large Language Models (LLMs) have shown strong reasoning capability in conversation, question answering and mathematics challenges. In this paper, we study if LLMs can be used to determine the equivalence between SQL queries under two notions of SQL equivalence (semantic equivalence and relaxed equivalence). To assist LLMs in generating high quality responses, we present two prompting techniques: Miniature & Mull and Explain & Compare. The former technique is used to evaluate the semantic equivalence in which it asks LLMs to execute a query on a simple database instance and then explore if a counterexample exists by modifying the database. The latter technique is used to evaluate the relaxed equivalence in which it asks LLMs to explain the queries and then compare if they contain significant logical differences. Our experiments demonstrate that using our techniques, LLMs are promising tools to help data engineers in writing semantically equivalent SQL queries, however challenges still persist, and have better alignment with human preference on text-to-SQL generations than execution accuracy.
@@ 1 Introduction
Structured Query Language (SQL) it is the standard programming language to query data systems and it is one of the most popular programming language Stackscale . Since SQL is not a procedural language, one may write different SQL queries that are semantically equivalent. As a result, determining the equivalence of SQL queries is a fundamental task in many practical important applications ranging from query optimizers in database systems Pirahesh et al. (1992); Bai et al. (2023); Zhou et al. (2021); Samwel et al. (2018) to evaluating the translation from user text-to-SQL Zelle and Mooney (1996); Zhong et al. (2020b); Khatry et al. (2023); Floratou et al. (2024a). However, it has been shown by the data management research community that checking the equivalence of two SQL queries is in general undecidable Abiteboul et al. (1995). While there are theoretical works to check equivalence under different assumptions Chandra and Merlin (1977); Sagiv and Yannakakis (1980); Cohen et al. (1999); Chu et al. (2018) and implemented prototypes such as Cosette Chu et al. (2017), they only support limited SQL operations and such projects are no longer maintained due to lack of experts. For instance, Cosette does not support commonly found operations like EXISTS, IN, queries with constant expressions, etc.; and Cosette is no longer maintained due to the lack of staff. Query optimizers and query processing engines in data systems still primarily rely on heuristic query rewriting rules Begoli et al. (2018). Exploring and writing these rules correctly require strong domain expertise and need many engineering hours.
@@
Moreover, due to the difficulty in evaluating the equivalence of two SQL queries, Text-to-SQL literature Zhong et al. (2017); Yu et al. (2018); Wang et al. (2019); Lee et al. (2021); Li et al. (2023) primarily use exact match (exact SQL string match) or execution match (compare the execution outputs) as the evaluation metrics. However, exact string match may overlook many equivalent SQL queries (there are many ways different ways to write the same SQL query), and execution match may misclassify different SQL queries as equivalent due to the lack of counterexamples in the underlining database instance. We note that there have been a few attempts that explore using semantic equivalence as the evaluation metrics Zhong et al. (2020a). These methods assess query semantic equivalence by generating numerous fuzzy database instances, leading to significant computational and storage demands. Specifically, evaluating a thousand pairs of SQL queries requires at least 75 minutes and consumes about 3 GB of storage space for creating these fuzz database instances Zhong et al. (2020a).
@@
Recent advancements in large language models (LLMs) have shown great capability in following instructions and in generating decisions that align with human judgement Bubeck et al. (2023); Chiang and Lee (2023); Yan et al. (2023a); Zhang et al. (2023b). For instance, It has been shown that GPT models are capable of performing first order logic reasoning  Saparov and He (2022), detecting implicit contents Huang et al. (2023); Zhang et al. (2023c), and solving complex logical problems Feng et al. (2023); Yao et al. (2023). In this paper, we investigate whether LLM can act as a judge to determine the equivalence of SQL queries. Hence, in this paper, we would like to mainly answer the following questions:
@@
  • Should LLMs be used to evaluate text-to-SQL generation?
@@ 2 Background
2.1 Motivations
We first introduce a query rewriting error that happened in the real-world system, known as the COUNT bug Ganski and Wong (1987). Here are two SQL queries Q1 and Q2 as shown below:
@@ 11 FROM Supply
One may think these two queries are equivalent, since they both return results based on the condition that qoh from Supply table equals to the counts of records in the Supply table when shipdate is less than 10; however they are different in how they handle the subquery. Suppose that there is a record $R$ in the Parts table such that R.qoh is zero and R.pnum does not exist in the Supply table. Then, Q1 will include record $R$ since R.qoh is zero and selecting the count for R.pnum from the Supply table will also return zero. On the other hand, Q2 will not include record $R$ because the TEMP result will not contain R.pnum.
Hence, real-world data systems need tools to avoid such pitfalls. While proving the semantic equivalency between two SQL queries has been proved to be undecidable Trahtenbrot (1963); Libkin (2004); Sagiv and Yannakakis (1980), we believe LLMs may offer new opportunities to help data engineers in scrutinizing their SQL optimization logic and potentially identify flaws accordingly. For instance, when we present the COUNT BUG to ChatGPT, ChatGPT is able to correctly identify that these two SQL queries are not equivalent.
Moreover, in text-to-SQL generation, many natural language questions are inherently ambiguous Katsogiannis-Meimarakis and Koutrika (2023); Floratou et al. (2024b), and as a result, they can be interpreted and answered in multiple ways Min et al. (2020). From the user’s perspective, different SQL queries that convey the same logical answer to a question may all be considered equivalent. However, the widely used execution match metric, which expects a single correct SQL query, does not capture these ambiguities. Consequently, we are interested in exploring whether LLMs can be utilized to assess the quality of text-to-SQL generation.
@@ 2.1.1 Two Notions of SQL Equivalence
Motivated from the query optimizers in data systems and evaluation metrics for text-to-SQL, we define two notions of SQL equivalence.
@@
2.2 Why LLM-SQL-Solver
First, LLM-SQL-Solver (semantic notion) can be directly used to find system bugs (e.g., COUNT BUG) and help engineers validate their SQL rewriting code bases. Second, since query optimizer relies on heuristic rules to find equivalent SQL forms, it may not explore all rewriting possibilities and hence, the rewritten query may lead to sub optimal performance. We believe that, with LLM-SQL-Solver, query optimizers can explore the vast universe of equivalent rewritten queries and then find the query plan with near optimal performance.
For text-to-SQL tasks, we believe that when comparing the gold SQL query (the correct answer) with the generated SQL query, the current most popular metric, execution match, does not reflects the human perspective and fails to capture a user’s opinion on the generated SQL. For instance, a common reasoning for failures in execution match is that the generated SQL selects fewer/extra columns than the gold SQL. In fact, a recent work has explicitly added "do not select extra columns" in its prompt to improve the execution accuracy Dong et al. (2023). However, from a user’s perspective providing extra columns may not reduce the quality of the answer. Also, there are many other similar reasons that can lead to false execution matches but not necessarily indicate a lower quality of the generated SQL query (such as the column order in selection, case sensitivity, duplicate entries, and columns in different tables share the similar contents). Moreover, execution matching fails to correctly identify logically different queries when the underlying database does not contain any counter examples, even when hundreds of human hours were consumed to create the benchmark Yu et al. (2018). Lastly, a fundamental limitation of execution accuracy is its need to query the underlying database. This is particularly pertinent as the text-to-SQL service provider usually does not have ownership of the user’s database and the user’s database may also contain sensitive information. As a result, it would be desirable to have a SQL solver determining the relax notion of equivalence without constructing sophisticated databases and without accessing any stored data. Hence, in additional to investigating the capability of foundational models in identifying semantic equivalence, we further explore whether LLM-SQL-Solver (relaxed notion) can be used as a better metric for text-to-SQL benchmarks.
@@ 3 Proposed Method
For semantic equivalence, we also added a short definition of SQL semantic equivalence (i.e., Note: Two SQL queries are semantically equivalent if and only if they yield identical results for all possible databases.). In addition, we first ask GPT to generate a counter example if there exists one and then decide if two SQL queries are equivalent. We call this prompt the Counter Example prompt. However, as shown in Figure 1, the counter example prompt leads to sub-optimal results, especially when GPT models only scratched the surface and observed the obvious (similarity between two SQL queries in structure and logic). This finding aligns with the prior works Wang et al. (2023); Zheng et al. (2023) which demonstrated that GPT models exhibit bias toward ignoring small differences when determining SQL equivalence and exhibit limited reasoning capability.
To overcome this observed phenomenon, we introduce our proposed prompting technique in detail to obtain better responses from large language models. In complex problem solving, people always start with a simple example and then introduce more complex elements to understand the full picture. Motivated by this simple intuition, we incorporate this idea and proposed the Miniature and Mull prompting technique. In particular, we first ask GPT to execute both queries on a simple database and then see if it can modify the database such that the two queries output different results (i.e., a counter example). As shown later in the experimental section, this approach can directly enhance the LLM’s response accuracy in determining SQL semantic equivalence. We believe that this prompting technique has wide applications in comparing the equivalence of two codes, plans, and logic.
For relaxed equivalence, we designed the Explain and Compare prompt by asking the LLMs to perform multi-step reasoning before determining if the two SQL queries are equivalent. This prompt follows the main ideas from the chain-of-thoughts Wei et al. (2022). We first prompt the LLMs to explain two SQL queries and then ask it to determine if these two queries contain significant difference in logic, as shown in Figure 1. We observe that LLMs are inclined to label two SQL queries as equivalent when they appear similar in structure. As a result, it is crucial for LLMs to be able to first spot the differences between these two queries and then apply this knowledge in considering if these two SQL queries are relaxed equivalent.
@@ 4 Evaluations
We evaluate the capabilities of popular large language models in identifying semantic and relaxed equivalence between a pair of SQL queries. In particular, we showcase the experimental results using the popular GPT 3.5 turbo (gpt-3.5-turbo) and GPT 4 turbo (gpt-4-1106-preview) models. The models are accessed via OpenAI API.
Hyper-parameter: We use zero demonstrations in our prompts. In addition, in the API call, we have set temperature to 0 which implies greedy sampling is used to generate the output; The max tokens are set to 3000; and the stopping tokens are set to "Answer:" for all experiments.
@@ 4.1 Data Source
First, we compare our proposed Miniature&Mull prompt with the CounterExample prompt on i) Non-equivalent Queries Chu et al. (2017) and ii) Equivalent Queries Zhong . These two data sources include pairs of SQL queries that have been expertly labeled as semantically non-equivalent or equivalent.
  • Nonequivalent Queries Chu et al. (2017): Four challenging and one easy questions, including the COUNT BUG, one exam question from CSE344 at the University of Washington, two bugs users found on SQL solver, and a simple hello world example.
  • Equivalent Queries Zhong et al. (2020a): A hundred pairs of expert labeled equivalent queries.
Next, we showcase results on the alignment of our proposed LLM-SQL-Solver judgements with human preferences on semantic and relaxed notions using the Spider benchmark Yu et al. (2018). The Spider benchmark contains 1034 development questions covering 20 databases in different domains. The Spider SQL queries are classified into four levels of complexity: easy, medium, hard, and extrahard (extra). We evaluate the judgments of LLM-SQL-Solver against execution accuracy (EX) based on human preferences. In all experiments, we use the open sourced EX evaluation code from GitHub  [^1] https://github.com/THU-BPM/ChatGPT-sql.
  • Chat Liu et al. (2023): Generate queries using ChatGPT with the zero-shot basic prompt.
  • DIN Pourreza and Rafiei (2023): Generate SQL queries by decomposing the task into fine grained sub-modules and use the GPT-4 model with few-shot prompting.
@@
4.2 Judging Semantic Equivalence
Judge Prompt Noneq Eq
GPT 35 turbo CounterExample .8 .45
Miniature&Mull 1.0 $\uparrow$ .57 $\uparrow$
GPT 4 turbo CounterExample .6 .7
Miniature&Mull 1.0 $\uparrow$ .83 $\uparrow$
Table 1: Results in determining semantic non-equivalent and equivalent queries.
Recall, two queries can not be semantically equivalent if there exists a counter example (a database instance) such that two queries produce different outputs. We used two different prompting methods Counter Example and Miniature & Mull to compare and contrast the accuracy of LLMs in determining semantic equivalence.
Judge Data Equivalence/Prompt Easy Medium Hard Extra Acc
@@ 4.2 Judging Semantic Equivalence
Table 2: A breakdown of the accuracy scores when using different metrics, i.e., EX and LLMs’ judgements with varying prompt techniques. The compared SQL queries are the gold SQL queries from Spider and the generated SQL queries from Chat, DIN, and DAIL.
As shown in Table 1, we can observe that the Miniature & Mull prompt always provides a higher accuracy than the Counter Example prompt. When using the Counter Example prompt (asking LLMs to determine SQL equivalence and provide a counter example if the SQL queries are not equivalent), GPT 4 turbo failed on a user found bug originally posted on Hackers News lukaseder . This bug contains two SQL queries where one used DISTINCT in selection and the other used UNION ALL to union two sub-queries. Hence, GPT 4 turbo failed to recognize that UNION ALL will not remove duplicate rows (though UNION will). GPT 3.5 turbo failed on the CSE344 exam question which involves different sub-queries. When using Miniature & Mull (asking LLMs to first run a simple example over the given SQL queries and then examining if the example database can be modified to generate a counter example), GPT 3.5 turbo and GPT 4 turbo both successfully identified all the mistakes. Our proposed Miniature & Mull prompt not only increased the accuracy of identifying non-equivalent SQL queries to 100% for both models but also increased the accuracy for identifying equivalent SQL queries by at least 12%. As a result, we believe that letting LLMs execute SQL queries on a simple database instance and then assessing whether a counterexample can be generated by modifying the database is a more effective method to determine SQL semantic equivalence.